OpenStudy (anonymous):

HELP ANYONE PLEASE? MEDAL+FAN! A $300 suit is on sale for 30% off. In addition to the sale, you have a 20% coupon that you found in your local paper. Determine the price that you will pay after the discounts. Round your answer to the nearest cent. a. $150.00 c. $168.00 b. $240.00 d. $210.00

OpenStudy (anonymous):

My first thought is A.

OpenStudy (asad786):

1 sec

OpenStudy (anonymous):

Ok :))

OpenStudy (asad786):

thats right

OpenStudy (asad786):

bacically 50% off

OpenStudy (welshfella):

300 - 20*300 ------ 100 then take 30% off

OpenStudy (asad786):

ok thats not right

OpenStudy (anonymous):

@welshfella how did you get that equation?

OpenStudy (welshfella):

= 240 then tale 30% off 240

OpenStudy (welshfella):

20% = 20 /100

OpenStudy (asad786):

welsh your wrong

OpenStudy (welshfella):

20% = 1/5 300 * 1/5 = 60 300 - 60 = 240

OpenStudy (anonymous):

By the way, the topic is about Consecutive Growth and Decay Factors.

OpenStudy (anonymous):

Why did you remove the 30% off? The questions saying is, A suit is 30% off and plus your coupon 20% off.

OpenStudy (welshfella):

240 * 30 --- = 72 100

OpenStudy (welshfella):

first I took 205 off then i will take 30% of the result

OpenStudy (welshfella):

* 20%

OpenStudy (anonymous):

Uggh, it's confusing dude.

OpenStudy (welshfella):

yea - I maybe wrong..... maybe you take 30% off 300 and 20% off 300 to be honest i'm confused also

OpenStudy (anonymous):

So the real answer is $150 RIGHT?

OpenStudy (welshfella):

this last method will give you 150 dollars whereas the method I used will give you 168 dollars

OpenStudy (welshfella):

I guess so

OpenStudy (anonymous):

BTW man, you did great to help me :))

OpenStudy (welshfella):

well - I try

OpenStudy (anonymous):

Thanks :)) @Asad786 @welshfella

OpenStudy (welshfella):

yw